# Quantile Advantage Estimation for Entropy-Safe Reasoning

### v1

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) strengthens LLM reasoning but training often oscillates between entropy collapse and entropy explosion. We trace both hazards to the mean-baseline used in value-free RL (e.g., GRPO & DAPO), which improperly penalizes negative-advantage samples under reward outliers. We propose Quantile Advantage Estimation (QAE), replacing the mean with a group-wise $K$-quantile baseline. QAE induces a response-level, two-regime gate: on hard queries ($p\leq 1{-}K$) it reinforces rare successes, while on easy queries (p>1Kp>1{-}K) it targets remaining failures. Under first-order softmax updates, we prove two-sided entropy safety, giving lower/upper bounds on one-step entropy change that curb explosion and prevent collapse. Empirically, this minimal modification stabilizes entropy, sparsifies credit assignment (with tuned $K$, roughly 80% of responses receive zero advantage), and yields sustained pass@1 gains on Qwen3-8B/14B-Base across AIME’24/’25 and AMC’23. These results identify baseline design—rather than token-level heuristics—as the primary mechanism for scaling RLVR  [^1] The code is available at https://github.com/junkangwu/QAE..
@@ 2 Preliminaries
GRPO further incorporates a KL divergence penalty against $\pi_{\bm{\theta}_{\text{old}}}$ to regularize the policy update.
@@ 3.2 An Analysis of Entropy Explosion in RLVR
In Figure 2, Clip-Higher triggers an early spike (steps 20–80) in anthropomorphic tokens—proposed by Yang et al. (2025b) as markers of “aha-moment” reasoning—that coincides with sharp pass@1 gains. However, after step 150, anthropomorphic token frequency returns toward baseline while performance plateaus. Thus, although Clip-Higher mitigates early collapse, its rapid escalation is coupled with entropy explosion that ultimately limits scaling.
@@ 5.1 Overall Performance across Models & Recipes
Qwen3-
8B-Base
Clip-Higher 32.71 56.66 39.69 71.23 92.11 97.50
@@
Qwen3-30B-
A3B-Base
GSPO 31.15 46.59 43.75 67.91 90.00 99.39
@@ 7 Related Work
RL has become a key technique for eliciting advanced reasoning in large language models (LLMs), a paradigm shift from its earlier applications in preference alignment via RLHF (Ouyang et al., 2022). This modern approach, termed Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert et al., 2024), leverages outcome-based optimization to achieve state-of-the-art performance in complex domains like mathematics and programming. Seminal works, including OpenAI’s o1 (ope, ) and DeepSeek R1 (DeepSeek-AI et al., 2025), demonstrated that RL can effectively scale reasoning capabilities, spurring a new line of research (Yang et al., 2025a; Team et al., 2025). Central to this progress are online, value-free algorithms that have generally outperformed offline preference optimization methods (Rafailov et al., 2023; Wu et al., 2024; 2025). In particular, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) and its successor, Dynamic Sampling Policy Optimization (DAPO) (Yu et al., 2025), have emerged as foundational baselines for many contemporary reasoning systems (Yue et al., 2025; Zeng et al., 2025; Hu et al., 2025). Our work uses DAPO as a representative algorithm to investigate a critical, unresolved challenge in this domain: the training instability caused by dysregulated policy entropy, which limits the performance and scalability of current RLVR methods.
@@
Evidence in RLVR links exploration tightly to entropy dynamics: performance gains concentrate on a minority of high-entropy “forking” tokens (Wang et al., 2025b), with “thinking tokens” acting as information peaks (Qian et al., 2025). At the sequence level, entropy can collapse early or explode if left unchecked (Cui et al., 2025). Extremes such as pure entropy minimization (Agarwal et al., 2025) or heavy upweighting of negative-advantage samples (Zhu et al., 2025) underscore the need for principled regulation, in line with cautions against indiscriminate maximum-entropy optimization (Zhang et al., 2025) and classic guidance to schedule target entropy (Xu et al., 2021) within regularized MDP theory (Geist et al., 2019; Ahmed et al., 2019). On the recipe side, entropy as advantage shaping (Cheng et al., 2025), Pass@k-based training (Chen et al., 2025), rubric-scaffolded exploration (Zhou et al., 2025), entropy-modulated policy gradients for long-horizon agents (Wang et al., 2025a), and outcome-based exploration (Song et al., 2025), together with agentic systems such as rStar2-Agent (Shang et al., 2025), provide practical means to prevent collapse/explosion while improving diversity.

### v2

# arXiv:2509.22611v2[cs.LG] 28 Feb 2026

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) strengthens LLM reasoning but training often oscillates between entropy collapse and entropy explosion. We trace both hazards to the mean-baseline used in value-free RL (e.g., GRPO & DAPO), which improperly penalizes negative-advantage samples under reward outliers. We propose Quantile Advantage Estimation (QAE), replacing the mean with a group-wise $K$-quantile baseline. QAE induces a response-level, two-regime gate: on hard queries ($p\leq 1{-}K$) it reinforces rare successes, while on easy queries (p>1Kp>1{-}K) it targets remaining failures. Under first-order softmax updates, we prove two-sided entropy safety, giving lower/upper bounds on one-step entropy change that curb explosion and prevent collapse. Empirically, this minimal modification stabilizes entropy, sparsifies credit assignment (with tuned $K$, roughly 80% of responses receive zero advantage), and yields sustained pass@1 gains on across AIME’24/’25 and AMC’23. These results identify baseline design—rather than token-level heuristics—as the primary mechanism for scaling RLVR  [^1] The code is available at https://github.com/junkangwu/QAE..
@@ 2 Preliminaries
GRPO further incorporates a KL divergence penalty against to regularize the policy update.
@@ 3.2 An Analysis of Entropy Explosion in RLVR
In Figure 2, Clip-Higher triggers an early spike (steps 20–80) in anthropomorphic tokens—proposed by Yang et al. (2025b) as markers of “aha-moment” reasoning—that coincides with sharp pass@1 gains. However, after step 150, anthropomorphic token frequency returns toward baseline while performance plateaus.
@@ 5.1 Overall Performance across Models & Recipes
Qwen3- 8B-Base Clip-Higher 32.71 56.66 39.69 71.23 92.11 97.50
@@
Qwen3-30B- A3B-Base GSPO 31.15 46.59 43.75 67.91 90.00 99.39
@@ 7 Related Work
RL has become a key technique for eliciting advanced reasoning in large language models (LLMs), a paradigm shift from its earlier applications in preference alignment via RLHF (Ouyang et al., 2022). This modern approach, termed Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert et al., 2024; Mroueh, 2025), leverages outcome-based optimization to achieve state-of-the-art performance in complex domains like mathematics and programming. Seminal works, including OpenAI’s o1 (13) and DeepSeek R1 (DeepSeek-AI et al., 2025), demonstrated that RL can effectively scale reasoning capabilities, spurring a new line of research (Yang et al., 2025a; Team et al., 2025). Central to this progress are online, value-free algorithms that have generally outperformed offline preference optimization methods (Rafailov et al., 2023; Wu et al., 2024; 2025). In particular, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) and its successor, Dynamic Sampling Policy Optimization (DAPO) (Yu et al., 2025), have emerged as foundational baselines for many contemporary reasoning systems (Yue et al., 2025; Zeng et al., 2025; Hu et al., 2025). Our work uses DAPO as a representative algorithm to investigate a critical, unresolved challenge in this domain: the training instability caused by dysregulated policy entropy, which limits the performance and scalability of current RLVR methods.
@@
Existing RLVR entropy research follows three strands: (i) Mechanistic analyses identifying where exploration concentrates, such as high-entropy “forking” tokens (Wang et al., 2025b) or “thinking tokens” (Qian et al., 2025), and the dynamics of sequence-level collapse or explosion (Cui et al., 2025); (ii) Objective-level regulation steering entropy via modified optimization targets (Agarwal et al., 2025; Zhu et al., 2025; Zhang et al., 2025) or regularized MDP scheduling (Geist et al., 2019; Ahmed et al., 2019; Xu et al., 2021); and (iii) Recipe/system-level heuristics injecting exploration via advantage shaping (Cheng et al., 2025), Pass@k training (Chen et al., 2025), agentic scaffolds (Zhou et al., 2025; Shang et al., 2025), and modulated gradients (Wang et al., 2025a; Song et al., 2025). Despite these advances, a baseline-level entropy control that is data-adaptive yet preserves standard policy-gradient updates remains missing. Our work fills this gap using a quantile baseline and binary masking, offering a drop-in lever that complements existing methods while explicitly targeting stability.
@@ Conclusion
Acknowledgments
This research is supported by the National Natural Science Foundation of China (U25A20445, 62572449, 62525211, 62302321). This research also benefited from the advanced computing resources provided by the Supercomputing Center of the USTC.